OpenStudy (anonymous):

Solve the following system of equations: −2x + y = 1 −4x + y = −1 (3, 1) (−1, 3) (−1, −3) (1, 3)

OpenStudy (tkhunny):

Go ahead and solve it. What's your plan?

OpenStudy (anonymous):

how do i solve it

OpenStudy (tkhunny):

Have you considered he methods in your book or that you discussed in class? Personally, I'm tempted to multiply the Top Equation by -1.

OpenStudy (anonymous):

should i try and isolate the Y

OpenStudy (tkhunny):

If you want. If not otherwise specified, you should feel free to use any desired method with which you are familiar.

OpenStudy (anonymous):

Y=3 X=1

OpenStudy (anonymous):

is that correct

OpenStudy (tkhunny):

Try it out and see. −2x + y = 1 −2(1) + (3) = 1 -2 + 3 = 1 -- Okay, that's the first one. −4x + y = −1 −4(1) + (3) = −1 -4 + 3 = -1 -- Okay, that's the second one. How did you arrive at that solution? Show your work!

OpenStudy (anonymous):

i used my scientific calculator

OpenStudy (tkhunny):

The old graphical method trick? That works, but you might learn less. Not necessarily.